OpenStudy (anonymous):

triangle ABC is congruent to triangle def. in triangle ABC,side AB measures 6, side BC measures 5x-21 and side CA measures 5. in triangle DEF side DE measures 6 side ER measures x-1 and side FD measures 5. what equation would help you to solve for the side length of BC and EF. please help

OpenStudy (anonymous):

please someone help

OpenStudy (anonymous):

as the triangles are congruent so their corresponding sides are equal now from analyzing the statement given AB=DE=6 ,CA=FD=5 and BC=EF BC=EF 5x-21=x-1 solve the equation

OpenStudy (anonymous):

okay so how do I solve it? im really confused right now

OpenStudy (anonymous):

don't freak out its easy 5x-21=x-1 rearrange the equation 5x-x=21-1 x=20/4 =5

OpenStudy (anonymous):

okay thank you so much for your help!! that made it a lot simpler!